6. Regardless of the GAAP depreciation method selected.. a. the maximum allowable depreciation over the assets life is the same. b. the total accumulated depreciation at the end of the assets life will equal the depreciable base. c. the total accumulated depreciation at the end of the assets life cannot exceed the depreciable base. d. all of the above. On August 1 of the current year, a company with a December 31 year-end buys a nonresidential building for $600,000, which includes land 7. that costs $100,000. Under MACRS, depreciation for this year will be a. $14,766 b. $12,305 c. $5,778 d. $4,815 8. For the building in Question 7, depreciation in the 12th year will be . a.$15,384 b. $12,820 c.$642 d.$535 9. When MACRS, Table 1 depreciation for an auto or light SUV, pickup or van is in excess of annual IRS limits the excess depreciation a. is lost permanently. b. may be taken by extending the MACRS recovery period for one extra year. c. may be taken by extending the MACRS recovery period for as many years as are needed to depreciate the cost basis. d. may result in the auto being depreciated under a GAAP method. During 2016, your firm acquires a 5,000-lb. SUV for $30,000 and a 6,200-lb. SUV for $35,000. How much of a Section 179 deduction will 10. your firm take on each vehicle in 2016? a. $25,000 and $25,000 b. $0 and $25,000 c. $30,000 and $35,000 d. You cannot take a Sec. 179 deduction for SUVs. 212
Answer #1

Answer :

6. d. All of the above

Regardless of the GAAP depreciation method selected the maximum allowable depreciation over the assets life will be same and total accumulated depreciation will be equal to depreciable base at end of asset's life. That is it cannot exceed depreciable base.

7. d. $ 4,815

The cost of non residential property = $ 600,000 - 100,000 = $ 500,000.

Rate of depreciation under Mid convention MACRS for 39 is 0.963% ( as the asset was purchased in 8th month- August of the current year ).

Depreciation = $ 500,000 × 0.963% = $ 4,815

8. b. $ 12,820

Rate of depreciation for years 2 to 39 is 2.564%

Depreciation for building in 12th year = 500,000 × 2.564 %= $ 12,820

9. c. May be taken by extending MACRS recovery period for as many years as are needed to depreciate the cost basis.

If the cost basis is not fully depreciated then, MACRS recovery period are extended as much needed to fully depreciate the asset.

10. b. $ 0 and $ 25,000

SUV 's with lbs greater than 6000 but within 14,000 pounds are allowed 179 deduction of $ 25,000.